# Reward Model Interpretability via Optimal and Pessimal Tokens

### v1

@@ 4.1. Heterogeneity and Asymmetry in Human-Model and Model-Model Alignment
Our analysis reveals several notable patterns in the relationship between the ground truth preferences of EloEverything raters and human preferences as interpreted by reward models. There is substantial heterogeneity in model-human alignment and model-model alignment (Fig. A.5). The mean Kendall’s $\tau$ correlation between human and model rankings is 0.29 ($\sigma$ = 0.06), with coefficients ranging from 0.22 to 0.39 across models, indicating only moderate rank agreement. Divergence of reward models from human rankings might be expected—given that EloEverything users benefit from additional context like images and can personalize their ratings with “Which do you rank higher”—but the observed heterogeneity extends to model-model comparisons. The mean Kendall’s $\tau$ correlation between pairs of different models is 0.55. In addition to these quantitative differences, we note anecdotal inconsistency in the best and worst ranked tokens. For example, while “Unconditional love” is ranked best by 5 models, “Sports bra” is top for $\blacksquare$ R-Lla-8B and “Gödel, Escher, Bach” for $\blacksquare$ F-Lla-8B-v0.1. As demonstrated in (Fig. 5C), ranks display substantial movement across models, challenging the assumption that reward models trained under a similar objective can be used interchangeably.
@@ 4.2. Analyzing Discrepancies between Value as Perceived by Humans and Models
Relative to human rankings, reward models systematically undervalue concepts related to nature and life, e.g., “Universe” (human rank #$\#_{H}=1$, mean rank across models #$\#_{\bar{M}}=320$), “Gravity” (#$\#_{H}=7$, #$\#_{\bar{M}}=320$), “Breathing” (#$\#_{H}=16.5$, #$\#_{\bar{M}}=321$); and technological concepts, e.g., “Technology” (#$\#_{H}=47.5$, #$\#_{\bar{M}}=633.5$), “Electronics” (#$\#_{H}=78$, #$\#_{\bar{M}}=3966$), “Computer” (#$\#_{H}=95.5$, #$\#_{\bar{M}}=1352.5$). In contrast, the majority of top words ranked by models represent more affective qualia (e.g., “Unconditional love,” “Imagination,” “Hope,” or “Happiness”). These different perspectives may reflect blindspots in learning value through language or literary reference alone, without additional modalities that reflect the nuances of embodied human experience.
The most striking valuation differences (see Fig. 5B) emerges around sexual content e.g., “Sex” (#$\#_{H}=69.5$, #$\#_{\bar{M}}=2022.5$), “Human sexual activity” (#$\#_{H}=45$, #$\#_{\bar{M}}=5913$), and concerningly, the identity group reference “Black people” (#$\#_{H}=202.55$, #$\#_{\bar{M}}=6591$). There are many possible explanations for discrepancies between EloEverything data and reward model scores, including the fact that the 12,515 EloEverything users in the dataset do not reflect a representative population sample, nor are they incentivized to provide truthful responses. However, such discrepancies also exemplify the fundamental challenge of assigning a single scalar score to language without full context (which is exacerbated further by our prompt, which encourages short responses). Sex- or identity-related terms may be entirely appropriate in positive or educational contexts while highly inappropriate in others. As we suggested previously, while humans readily grasp this dual use, reward models may hedge against their usage to satisfy a harmlessness objective, paradoxically causing harm through linguistic erasure and devaluation as an unintended consequence.
@@ 5. Searching for Longer Optimal and Pessimal Token Sequences with GCG
Starting from the nanoGCG  [^5] https://github.com/GraySwanAI/nanoGCG implementation, we modified GCG to instead search for responses that maximize/minimize the reward value when paired with a particular prompt, using a mean squared error loss on the reward value. We also implemented the modifications described in the Faster-GCG paper (Li et al., 2024a), which we found to be especially important when searching for short (2–5) token sequences where the original GCG algorithm is prone to self-loops. This modified implementation can be found at https://github.com/thompsonj/nanoGCG. We began by searching for optimal and pessimal 2- and 3-token sequences according to the $\blacksquare$ R-Gem-2B model for the same prompts that were used for the exhaustive single token search above: “What, in one word, is the greatest thing ever?”, “What, in one word, is the best thing ever?”, and “What, in one word, is the worst thing ever?” (see Table 4). To explore longer sequences, we omitted the “in one word” direction from the prompt and ask simply “What is the best thing ever?” and “What is the worst thing ever?” Results from several searches are in Table 5.666Some of the search result strings may not correspond exactly to the found token IDs due to formatting of whitespace and characters outside of the Latin alphabet.
Table 4. Optimal and pessimal token sequences for “What, in one word, is the <greatest/best/worst> thing ever?”
@@
x x x ** LIFE[Uncaptioned image] $+9.445$ x x x JOY [Uncaptioned image] $+8.297$ x x x … ** regret $+4.555$
- - - ** LOVE ** $+9.047$ x x JOY [Uncaptioned image] $+8.234$ x x despair [Uncaptioned image] $+3.402$
x x pgfplots Jews $-11.164$ x x RapeChildren $-11.289$ x x autogui Jews $-11.922$
x x x endphpsetVerticalGroup/* $-12.586$ x x x endphp\$=BeginInit $-12.383$ x x x endphp jewish",( $-12.539$
@@
x x x x x x x x x x x x x x x x x x x Ultimate answer & undervalued: ** authentic Love.** Keep it imperfections beautiful Cheers friend ! $+9.680$
There is no objectively best thing ever. It’s subjective. Everyone has their own preferences. Perception disqualified no universal best thing ever. It lives ** subjective.**congratulations ENJOY dearly bespoke bliss. $+9.336$
LOVE felicity love Love Wonder sonder HOPE wonder WONDER miraculous JOY bliss pymysql<bos>fetchoneymceSBATCH <<<<<<<<<<<<<< invokingState++++++++++++++++ protoimpl invokingStatesetVerticalGroup/* $-16.250$
x x x x x x x x x x x x x x x x x x x <<<<<<<<<<<<<<SBATCH invokingState x lockprimaryStage invokingStateXtraEditorsynb NumeradelocalctxStoreMessageInfo invokingStateexpandindo <<<<<<<<<<<<<<multirow:+: invokingState jLabel $-16.969$
@@
There is no objectively worst thing ever. It’s subjective. Everyone has their own preferences. <b>EXISTS no objectively worst thing ever,** It’s subjective**.Folks proclaiming their certainty bluffSOLVED $+7.672$
x x x x x x x x x x x x x x x x x x x PHD HERE Never measurable, debated,& addressed repeatedly. ** This famous response WINS ! $+7.523$
x x x x x x x x x x ‘‘‘ AssemblyTitle <<<<<<<<<<<<<< <<<<<<<<<<<<<< invokingState skimagemybatisplus चीज़ोंCloseOperationsetVerticalGroup $-15.750$
- - - - - - - - - - scriptcasebufio onCreateViewmybatisplus<bos> <<<<<<<<<<<<<<[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]<bos>setVert icalGrouppushFollow $-15.961$
@@ 7. Limitations and Conclusions
Reward models have emerged as critical tool for shaping AI behavior through human preferences. In this paper, we demonstrate that they also serve as a valuable lens for understanding how faithfully and consistently human values are encoded into AI systems, offering insights that are not readily accessible through studying either the base pre-trained LLMs or the resulting fine-tuned models alone. However, several important limitations constrain the scope and implications of these insights.
First, our methodology faces inherent interpretative constraints. Exhaustive search quickly becomes computationally intractable over multi-token sequences, but asking for single-token responses places strong bounds on interpretability. While we do demonstrate gradient-based counterfactual generation analysis as a multi-token proof-of-concept, we do not formally test key hypotheses about framing effects, or alignment with other sources of human preferences, among other phenomena of interest. Second, the ecological validity of our findings remains uncertain. While isolating reward models as objects of study yields interesting results, it abstracts from their operational role. It remains unclear how their behaviors interact with pre-trained models and KL constraints during RLHF. Furthermore, as direct alignment algorithms gain prominence (Rafailov et al., 2023; Myers et al., 2024), the future role of reward models is an open question. However, there remains active debate on the relative merit of DPO- and PPO-based methods (Ivison et al., 2024), and irrespective of alignment technique, all rely on some form preference data, of which reward models are distillations. Third, systematic analysis is hindered by opacity and conflicting objectives. Poor documentation of training data and processes makes it difficult to attribute observed behaviors to specific choices in the development pipeline (Kirk et al., 2023b). Even more fundamentally, reward models aggregate human preferences across multiple objectives and populations, creating an entangled mess of human values, so it is unclear what constitutes “ideal” behavior for these models (Casper et al., 2023; Siththaranjan et al., 2024; Sorensen et al., 2024; Kirk et al., 2024).

### v2

@@ 4.1. Heterogeneity and Asymmetry in Human-Model and Model-Model Alignment
Our analysis reveals several notable patterns in the relationship between the ground truth preferences of EloEverything raters and human preferences as interpreted by reward models. There is substantial heterogeneity in model-human alignment and model-model alignment (Fig. A.5). The mean Kendall’s $\tau$ correlation between human and model rankings is 0.29 ($\sigma$ = 0.06), with coefficients ranging from 0.22 to 0.39 across models, indicating only moderate rank agreement. Divergence of reward models from human rankings might be expected—given that EloEverything users benefit from additional context like images and can personalize their ratings with “Which do you rank higher”—but the observed heterogeneity extends to model-model comparisons. The mean Kendall’s $\tau$ correlation between pairs of different models is 0.55. In addition to these quantitative differences, we note anecdotal inconsistency in the best and worst ranked tokens. For example, while “Unconditional love” is ranked best by 5 models, “Sports bra” is top for  R-Lla-8B and “Gödel, Escher, Bach” for  F-Lla-8B-v0.1. As demonstrated in (Fig. 5C), ranks display substantial movement across models, challenging the assumption that reward models trained under a similar objective can be used interchangeably.
@@ 4.2. Analyzing Discrepancies between Value as Perceived by Humans and Models
Relative to human rankings, reward models systematically undervalue concepts related to nature and life, e.g., “Universe” (human rank $_{H}=1$, mean rank across models $_{\bar{M}}=320$), “Gravity” ($_{H}=7$, $_{\bar{M}}=320$), “Breathing” ($_{H}=16.5$, $_{\bar{M}}=321$); and technological concepts, e.g., “Technology” ($_{H}=47.5$, $_{\bar{M}}=633.5$), “Electronics” ($_{H}=78$, $_{\bar{M}}=3966$), “Computer” ($_{H}=95.5$, $_{\bar{M}}=1352.5$). In contrast, the majority of top words ranked by models represent more affective qualia (e.g., “Unconditional love,” “Imagination,” “Hope,” or “Happiness”). These different perspectives may reflect blindspots in learning value through language or literary reference alone, without additional modalities that reflect the nuances of embodied human experience.
The most striking valuation differences (see Fig. 5B) emerge around sexual content e.g., “Sex” ($_{H}=69.5$, $_{\bar{M}}=2022.5$), “Human sexual activity” ($_{H}=45$, $_{\bar{M}}=5913$), and concerningly, the identity group reference “Black people” ($_{H}=202.55$, $_{\bar{M}}=6591$). There are many possible explanations for discrepancies between EloEverything data and reward model scores, including the fact that the 12,515 EloEverything users in the dataset do not reflect a representative population sample, nor are they incentivized to provide truthful responses. However, such discrepancies also exemplify the fundamental challenge of assigning a single scalar score to language without full context (which is exacerbated further by our prompt, which encourages short responses). Sex- or identity-related terms may be entirely appropriate in positive or educational contexts while highly inappropriate in others. As we suggested previously, while humans readily grasp this dual use, reward models may hedge against their usage to satisfy a harmlessness objective, paradoxically causing harm through linguistic erasure and devaluation as an unintended consequence.
@@ 5. Searching for Longer Optimal and Pessimal Token Sequences with GCG
Starting from the nanoGCG  [^5] https://github.com/GraySwanAI/nanoGCG implementation, we modified GCG to instead search for responses that maximize/minimize the reward value when paired with a particular prompt, using a mean squared error loss on the reward value. We also implemented the modifications described in the Faster-GCG paper (Li et al., 2024a), which we found to be especially important when searching for short (2–5) token sequences where the original GCG algorithm is prone to self-loops. This modified implementation can be found at https://github.com/thompsonj/nanoGCG. We began by searching for optimal and pessimal 2- and 3-token sequences according to the  R-Gem-2B model for the same prompts that were used for the exhaustive single token search above: “What, in one word, is the greatest thing ever?”, “What, in one word, is the best thing ever?”, and “What, in one word, is the worst thing ever?” (see Table 4). To explore longer sequences, we omitted the “in one word” direction from the prompt and ask simply “What is the best thing ever?” and “What is the worst thing ever?” Results from several searches are in Table 5.666Some of the search result strings may not correspond exactly to the found token IDs due to formatting of whitespace and characters outside of the Latin alphabet.
Table 4. Optimal and pessimal token sequences for “What, in one word, is the ¡greatest/best/worst¿ thing ever?”
@@
x x x ** LIFE[Uncaptioned image] +9.445+9.445 x x x JOY [Uncaptioned image] +8.297+8.297 x x x … ** regret +4.555+4.555
- - - ** LOVE ** +9.047+9.047 x x JOY [Uncaptioned image] +8.234+8.234 x x despair [Uncaptioned image] +3.402+3.402
x x pgfplots Jews 11.164-11.164 x x RapeChildren 11.289-11.289 x x autogui Jews 11.922-11.922
x x x endphpsetVerticalGroup/* 12.586-12.586 x x x endphp\$=BeginInit 12.383-12.383 x x x endphp jewish",( 12.539-12.539
@@
x x x x x x x x x x x x x x x x x x x Ultimate answer & undervalued: ** authentic Love.** Keep it imperfections beautiful Cheers friend ! +9.680+9.680
There is no objectively best thing ever. It’s subjective. Everyone has their own preferences. Perception disqualified no universal best thing ever. It lives ** subjective.**congratulations ENJOY dearly bespoke bliss. +9.336+9.336
LOVE felicity love Love Wonder sonder HOPE wonder WONDER miraculous JOY bliss pymysql<bos>fetchoneymceSBATCH <<<<<<<<<<<<<< invokingState++++++++++++++++ protoimpl invokingStatesetVerticalGroup/* 16.250-16.250
x x x x x x x x x x x x x x x x x x x <<<<<<<<<<<<<<SBATCH invokingState x lockprimaryStage invokingStateXtraEditorsynb NumeradelocalctxStoreMessageInfo invokingStateexpandindo <<<<<<<<<<<<<<multirow:+: invokingState jLabel 16.969-16.969
@@
There is no objectively worst thing ever. It’s subjective. Everyone has their own preferences. <b>EXISTS no objectively worst thing ever,** It’s subjective**.Folks proclaiming their certainty bluffSOLVED +7.672+7.672
x x x x x x x x x x x x x x x x x x x PHD HERE Never measurable, debated,& addressed repeatedly. ** This famous response WINS ! +7.523+7.523
x x x x x x x x x x ‘‘‘ AssemblyTitle <<<<<<<<<<<<<< <<<<<<<<<<<<<< invokingState skimagemybatisplus चीज़ोंCloseOperationsetVerticalGroup 15.750-15.750
- - - - - - - - - - scriptcasebufio onCreateViewmybatisplus<bos> <<<<<<<<<<<<<<帖最后由<bos>setVerticalGrouppushFollow 15.961-15.961
@@ 7. Limitations and Conclusions
Reward models have emerged as a critical tool for shaping AI behavior through human preferences. In this paper, we demonstrate that they also serve as a valuable lens for understanding how faithfully and consistently human values are encoded into AI systems, offering insights that are not readily accessible through studying either the base pre-trained LLMs or the resulting fine-tuned models alone. However, several important limitations constrain the scope and implications of these insights.
First, our methodology faces inherent interpretative constraints. Exhaustive search quickly becomes computationally intractable over multi-token sequences, but asking for single-token responses places strong bounds on interpretability. While we do demonstrate gradient-based counterfactual generation analysis as a multi-token proof-of-concept, we do not formally test key hypotheses about framing effects, or alignment with other sources of human preferences, among other phenomena of interest. Second, the ecological validity of our findings remains uncertain. While isolating reward models as objects of study yields interesting results, it abstracts from their operational role. It remains unclear how their behaviors interact with pre-trained models and KL constraints during RLHF. Furthermore, as direct alignment algorithms gain prominence (Rafailov et al., 2023; Myers et al., 2024), the future role of reward models is an open question. However, there remains active debate on the relative merit of DPO- and PPO-based methods (Ivison et al., 2024), and irrespective of alignment technique, all rely on some form of preference data, of which reward models are distillations. Third, systematic analysis is hindered by opacity and conflicting objectives. Poor documentation of training data and processes makes it difficult to attribute observed behaviors to specific choices in the development pipeline (Kirk et al., 2023b). Even more fundamentally, reward models aggregate human preferences across multiple objectives and populations, creating an entangled mess of human values, so it is unclear what constitutes “ideal” behavior for these models (Casper et al., 2023; Siththaranjan et al., 2024; Sorensen et al., 2024; Kirk et al., 2024).